OpenStudy (anonymous):

If a+b=5 and x+y=3, what is −2x−5a−5b−2y?

OpenStudy (loser66):

what do you think?

OpenStudy (phi):

**−2x−5a−5b−2y?*** try writing it like this: -2x + -2y + -5a + -5b now "factor out" -2 from the first two terms also, factor out -5 from the other two terms what do you get ?

OpenStudy (phi):

refresh your browser if the post does not display properly

OpenStudy (anonymous):

\[-2y+-5b\] ?

OpenStudy (anonymous):

-31

OpenStudy (phi):

when you factor, the x (and a) do not disappear Here is an example: 2(3+4) means 2*3 + 2*4 (you can check: 2*(3+4) is 2*7= 14, and this matches 2*3=6 added to 2*4= 8: 6+8= 14.) anyway: if you see 2*3 + 2*4 you can "undo the distribute" by factoring out the 2 you get 2*(3+4) if you have letters it works the same way example: -2a + -2b you have -2 in both terms. factor it out: -2(a+b)